OpenStudy (anonymous):

Identify the range of the equation y = −x2 − 8x − 10. All real numbers y ≥ −2 y ≤ 6 y ≤ −4 @cwrw238

OpenStudy (anonymous):

is it suppose to be -x^2

OpenStudy (anonymous):

it's suppose to be the range not the solution.

OpenStudy (debbieg):

is it: \(y=-x^2-8x-10\)?

OpenStudy (debbieg):

I'm assuming so - so what do you know about the range of a quadratic?

OpenStudy (anonymous):

yes it is and that is my mistake I very sorry. This is a pre-assessment to a new module.

OpenStudy (debbieg):

OK, so what do you know about the range of a quadratic function?

OpenStudy (anonymous):

I don't have a clue to what it is or even how or work it. :D

OpenStudy (debbieg):

What do you know about quadratic functions? (e.g., equations of the form \(y=ax^2bx+c\)? Do you know what shape the graph of such a function is? Do you know what the range of a function is?

OpenStudy (debbieg):

oops, sorry that should be \[y=ax^2+bx+c\]up there.

OpenStudy (anonymous):

I know its a parabola and it open downwards, Correct?

OpenStudy (debbieg):

Yes, good! Now, do you know how to find the vertex?

OpenStudy (anonymous):

I guess that the range is inside the side that the parabola opens.

OpenStudy (debbieg):

Hmmmm.... no, the range of a function (in general) is all of the possible y values that the function can take on. In terms of a graph, you can "see" the range as all the values of y that have a point on the graph associated to them.

OpenStudy (debbieg):

So, e.g., if the function was y=x^2, the most basic quadratic, it would have its vertex at (0,0) and open up. So for every point on the graph, \(y\ge0\) so that is the range.

OpenStudy (anonymous):

Ok well how do I find the vertex?

OpenStudy (debbieg):

So for a parabola, if you know the vertex and direction of opening, you know the range - see? Because the y of the vertex is either the lower bound of the range (if it opens upward) or the upper bound (if it opens downward).

OpenStudy (anonymous):

So since it opens downwards it would be less than or equal too?

OpenStudy (debbieg):

There are several different ways. :) If all you need is the range, I would probably use the formula \(\large x=\dfrac{-b}{2a}\) to find the x-coordinate of the vertex. Once you have that, plug it back into the equation, and that will give you the y coordinate of the vertex.

OpenStudy (debbieg):

Yes, since it opens down, it will be a less than expression. :)

OpenStudy (anonymous):

Ok which one is b and which one is a?

OpenStudy (anonymous):

@cwrw238

OpenStudy (debbieg):

Like I said above, the equation is of the form \(y=ax^2+bx+c\). So you tell me: what is a, what is b?